Solve using the substitution method:
x-y=5
x+2y=7
and please explain...thank you

Answer by jim_thompson5910(28550)

Isolate x in the first equation x-y=5 to get

x-y=5

x=5+y

Now plug this into the second equation

x+2y = 7

5+y+2y = 7 ... replace 'x' with 5+y (since x=5+y)

*** Notice how this completely gets rid of 'x' in the second equation. This leaves us with y.

Now solve for y

5+y+2y = 7

5+3y = 7

3y = 7-5

3y = 2

y = 2/3

Now that we know that y = 2/3, we can use this to find x.

x = 5+y

x = 5+2/3

x = 15/3+2/3

x = 17/3

So the solutions are x = 17/3 and y = 2/3

These solutions can be represented as the ordered pair (17/3, 2/3)